P5.6 Estimate the time it takes send a file that has 20000 characters using RS232-serial interfacing if the baud rate is set at 38400 bps using a 10-bit data packet.
Answer #1

Hi,

Assuming that each character is 8-bit long, plus one start bit and one stop bit (thus making it a total of 10-bits in a packet), the time taken is computed as:

Total_Time_taken = Time_per_bit * Number_of_bits_per_packet*Number_of_characters

Time_per_bit = (1/38400) sec

Therefore Total_Time_Taken = (1/38400)*10*20000 = 5.2083 seconds

Hope this helps